OpenStudy (anonymous):

I'm having a hard time with "percent of". If you could please tell how to solve them and give an explanation on how you solved it (which I think is the same exact thing), I would be much appreciated. Here's are the questions. 1.What percent of 20 is 30? 2.What percent of 20 is 30? 3.45% of what is 9?

OpenStudy (anonymous):

hi @paki

OpenStudy (paki):

hey

OpenStudy (anonymous):

whats going on

OpenStudy (tkhunny):

What percent of 100 is 40? ==> 40/100 = 0.40 = 40% What percent of 100 is 80? ==> 80/100 = 0.80 = 80% What percent of 100 is 100? ==> 100/100 = 1.00 = 100% What percent of 100 is 175? ==> 175/100 = 1.75 = 175% What percent of 100 is 3? ==> 3/100 = 0.03 = 3% Same thing. Every time.

OpenStudy (anonymous):

@tkhunny thank you. I will remember that :)